# Sinaĭ excursions

### v4

@@ 1.1. Sinaĭ walks
As discussed by Aurzada and Simon [2, Section 3], in their survey on persistence probabilities, Sinaĭ’s proof is based on the sequence of times $0=\tau_{0},\tau_{1}\ldots$ that the walk visits 0. This gives rise to another random walk, whose increments are the signed areas accumulated between these times. A key ingredient is Sparre Andersen’s [19] classical result that, for $|x|\leqslant 1$,
@@
See also [2, Section 4] for more on Sinaĭ’s [18] motivation, in relation to the inviscid Burgers’ equation. Roughly speaking, this equation models a turbulent fluid (with no viscosity) which gives rise to “shock waves.” When the system is started with self-similar, Brownian data, the exponent $1/4$ in (1) is related to the fact that the set of initial positions of particles not yet “shocked” by time $t=1$ has Hausdorff dimension $1/2$; see [2, 18, 7, 13].
@@ 1.2. Sinaĭ excursions
We call $(S_{0},S_{1},\ldots,S_{4n})$ a Sinaĭ excursion if it is a Sinaĭ walk and $S_{4n}=A_{4n}=0$. Such renewal times are only possible at multiples of $4$.
Sinaĭ excursions play a role in the work of Caravenna and Deuschel [8]. In this context, they are random polymer chains exhibiting entropic repulsion, related to the interface in the wetting model that forms between a gas pressed diffusively against a surface by a liquid; see [8, 2, 20] for more details. We also note that Sinaĭ excursions are the discrete analogue of the positive Kolmogorov excursions (from zero and back) studied recently by Bär, Duraj and Wachtel [5].
@@
verifying a conjecture in [8].
@@
To give a first hint about the connection between $p_{n}$ and $\Xi_{n}$, consider the $2n$ times $t_{1}<\cdots<t_{2n}$ before down steps in a walk $(S_{0},S_{1},\ldots,S_{4n})$ when $S_{t+1}-S_{t}=-1$. By Lemma 3.1 below, if $A_{4n}=0$ then $\sum_{j=1}^{2n}t_{j}=n(4n-1)$, with is equal to $3n$ mod $4n$. More to the point, as we will see, $\Xi_{n}$ is related to the number of bridges that can be turned into a Sinaĭ excursion, by cyclically shifting their increments.
@@ 1.3. A Sparre Andersen analogue
$\varphi_{n}={\bf P}(A_{1},\ldots,A_{4n}\geqslant 0,\>A_{4n}=Y_{4n}=0).$ (6)
By the local limit in [1, Proposition 2.1], it follows that
$n^{2}{\bf P}(A_{4n}=Y_{4n}=0)\to\frac{\sqrt{3}}{4\pi}.$
@@ 1.4. Transferring asymptotics
A probability distribution $(q_{n},n\geqslant 0)$ is sub-exponential if $q_{n}/q_{n+1}\to 1$ and $q_{n}^{*}/q_{n}\to 2$, where
@@
In [11] it is proved that if $p_{n}$ and $\nu_{n}$ are related by (10), then $\nu_{n}/\lambda$ is sub-exponential if and only if $p_{n}\sim\nu_{n}$ and $\nu_{n}/\nu_{n+1}\to 1$. Intuitively, this follows by the “one big jump principle.” A large value of $L_{1}$ is likely due to one big jump by time $t=1$. As a special case, Hawkes and Jenkins [14] showed that $p_{n}\sim\nu_{n}$ if $\nu_{n}$ is regularly varying with index $\gamma<-1$.
@@ 1.5. Outline
We have shown how Theorem 1.1 follows by Theorem 1.2. The remainder of the article is devoted to the proof of Theorem 1.2.
@@
Next, we claim that, for integers $10\leqslant m_{1}\leqslant\cdots\leqslant m_{2n}\leqslant 2n-$ 1, we have that $\sum_{i=1}^{2n}m_{i}$ is equal to 00 mod $2n$ if and only if $\sum_{i=1}^{2n}(m_{i}+i)$ is equal to $n$ or $3n$ mod $4n$. Indeed, to see this, simply note that $\sum_{i=1}^{2n}i=n(2n+1)$ is equal to $n$ or $3n$ mod $4n$ (depending on the parity of $n$). ∎
@@ 3. Times before down steps
denote the sequences of times before down steps, that is, times $10\leqslant t\leqslant 2n-$ 1 such that $\Delta B_{t}=-1$.
@@
$\sum_{k=1}^{4n}B_{k}=\sum_{k=0}^{4n-1}(4n-k)\Delta B_{k}=\sum_{k=0}^{4n-1}(4n-k)-2\sum_{j=1}^{n}(4n-t_{j}),$
which simplifies to $-2n(4n-1)+2\sum_{i=1}^{2n}t_{i}$, as claimed. ∎
@@ 4. Sparre Andersen for Sinaĭ excursions
The key to proving (13) is to observe that any Sinaĭ excursion can be decomposed into a series of irreducible Sinaĭ excursions. More specifically, let $\Phi_{n}^{+}$ be the number of walks $(S_{0},S_{1},\ldots,S_{4n})$ for which $S_{0}=A_{0}=0$, $S_{4n}=A_{4n}=0$ and $A_{1},\ldots,A_{4n-1}>0$. Then $\Phi_{n}$ is a renewal sequence, in the sense that
$\Phi(x)=\sum_{m=0}^{\infty}[\Phi^{+}(x)]^{m}=\frac{1}{1-\Phi^{+}(x)},$
where $\Phi(x)=\sum_{n}\Phi_{n}x^{n}$ and $\Phi(x)=\sum_{n}\Phi_{n}^{+}x^{n}$. In other words, each Sinaĭ excursion is a series of $m$ many irreducible Sinaĭ excursions, for some $m\geqslant 1$.
Figure 2. Marked objects of size $n$ ($nA_{n}$ many) can be split into an unmarked object of size $n-k$ ($A_{n-k}$ many) and an object of size $k$ with a mark in its first irreducible part ($A_{k}^{*}$ many). Vertical lines delimit irreducible parts. Recurrence (14) follows, summing over $k$.
@@
where $A_{n}^{\prime}$ is the number of pairs $(X,s)$, where $X\in\mathcal{A}_{n}$ and $0\leqslant m<\ell$, where $\ell$ is the size of its first irreducible part (i.e., its first part is in $\mathcal{A}_{\ell}$). In other words, $A_{n}^{\prime}$ is the number of marked objects of size $n$, where the mark is in their first irreducible part.
This fact is proved by Bassan and the authors in [6], using the following simple observation. Note that $nA_{n}$ counts the number of marked objects of size $n$, i.e., pairs $(X,i)$ with $X\in\mathcal{A}_{n}$ and $1\leqslant i\leqslant n$. On the other hand, consider the sub-object consisting of the part containing $i$ and all subsequent parts. This object is of size $k$, for some $k\geqslant 1$, and is marked somewhere in its first irreducible part. All other previous parts form some unmarked object of size $n-k$. Hence, there are equivalently $\sum_{k=1}^{n}A_{k}^{\prime}A_{n-k}$ many such pairs $(X,i)$. See Figure 2.
To complete the proof of Theorem 1.2, we show the following.
@@
  • from the set of all pairs $({\mathcal{B}},j)$, where ${\mathcal{B}}$ is a Sinaĭ excursion of length $4n$, with first positive Sinaĭ excursion of length $\ell=4k$, and $1\leqslant j\leqslant 2k$ is an integer,
@@
To describe $\Upsilon$, consider such a ${\mathcal{B}}$, as above. Let
@@
be the times before up steps in the first positive Sinaĭ excursion of ${\mathcal{B}}$. Let ${\mathcal{B}}^{(j)}$ be the bridge obtained from ${\mathcal{B}}$ by cyclically shifting ${\mathcal{B}}$ to the left by $i_{j}$. In other words, the $k$th increment of ${\mathcal{B}}^{(j)}$ is the $(k+i_{j})$th (understood mod $4n$) increment of ${\mathcal{B}}$. In particular, ${\mathcal{B}}^{(1)}={\mathcal{B}}$. We let $\Upsilon({\mathcal{B}},j)$ to be the set of times before down steps in ${\mathcal{B}}^{(j)}$. Since ${\mathcal{B}}$ is a Sinaĭ excursion its times before down steps sum to $3n$ mod $4n$, by Lemma 3.1. We shift all of these $2n$ times by the same amount to obtain the times before down steps in ${\mathcal{B}}^{(j)}$, so these sum to $n$ or $3n$ mod $4n$. See Figure 3 for an example.
@@
5. Final remarks
We note that Vysotsky [21] has sharpened and generalized Sinaĭ’s persistence probability (1), showing that
${\bf P}(A_{1},\ldots,A_{n}\geqslant 0)\sim Cn^{-1/4},$
Finally, let us mention that, as a consequence of considerably more technical arguments than those in the current article, we [10, Corollary 4] recently proved that
@@ 5. Final remarks
where $\tau=\inf\{t:Y_{t}=0,\>A_{t}\leqslant 0\}$. Informally, $\tau$ is the first time that a random walk is “at risk” of not being Sinaĭ. Our main Theorem 1.1 above implies that ${\bf P}(A_{\tau}=0)=1-e^{-\lambda}$, where $\lambda$ is as in (8), so that
@@ 6. Acknowledgments
SD acknowledges the financial support of the CogniGron research center and the Ubbo Emmius Funds (University of Groningen). BK was partially supported by a Florence Nightingale Bicentennial Fellowship (Oxford Statistics) and a Senior Demyship (Magdalen College). BK thanks Vasu Tewari for indicating the work of von Sterneck.

### v5

@@ 1. Introduction
1.1. Persistence probabilities
There is a rich history of calculating persistence probabilities in mathematics, wherein we ask for a random process to continue to satisfy some property of interest. For instance, Bertrand’s ballot theorem [1] from 1887 can be viewed as an early example. No less famous is Sparre Andersen’s [28] formula from 1954, concerning the probability that a simple random walk remains positive.
Applications of persistence in statistical physics began in the late 1980s, as it relates to, e.g., fluctuating interfaces and sticky particle systems. We refer to the surveys [23, 13, 10, 3] for a detailed overview of existing results and applications.
In this work, we introduce an analogue of Sparre Andersen’s formula for the area process of a random walk; see (2) and Theorem 1.2 below.
Persistence probabilities for such area processes were first studied by Sinaĭ [26], in relation to the inviscid Burgers’ equation. This equation models a turbulent fluid which gives rise to shock waves. Such discontinuities are overcome mathematically via control surfaces, i.e., interfaces through which material flows from one side of a discontinuity to the other. When the system is started with self-similar, Brownian data, the exponent $1/4$ in (1) below is related to the fact that the set of initial positions of particles not yet “shocked” by time $t=1$ has Hausdorff dimension $1/2$; see [3, 27, 8, 21].
Times when the random walk and its area process revisit 0 are renewal times for such persistence problems. We call the trajectories between such times Sinaĭ excursions; see Section 1.3 below. These excursions play a key role in the work of Caravenna and Deuschel [11] on the wetting model. In this context, Sinaĭ excursions are related to random polymer chains exhibiting a phenomenon referred to as entropic repulsion. More specifically, the area process of a Sinaĭ excursion is used to model the interface that forms between a gas pressed diffusively against a surface by a liquid; see [11, 3, 9, 29, 20] for more details.
Our main result Theorem 1.1 (see also Corollary 5.1) identifies the precise asymptototic probability that a random walk is a Sinaĭ excursion. This sharpens the original bounds (up to polylogarithmic factors) by Caravenna and Deuschel [11] and the subsequent improvement (up to constant factors) by Aurzada, Dereich and Lifshits[2]. Our proof utilizes our area analogue of Sparre Andersen’s formula mentioned above, old formulas from additive number theory proved in the early 1900s by von Sterneck [4], and the Tauberian theorems related to the Lévy–Khintchine formula proved by Hawkes and Jenkins [22].
@@ 1.2. Sinaĭ walks
As discussed by Aurzada and Simon [3, Section 3], in their survey on persistence probabilities, Sinaĭ’s proof is based on the sequence of times $0=\tau_{0},\tau_{1}\ldots$ that the walk visits 0. This gives rise to another random walk, whose increments are the signed areas accumulated between these times. A key ingredient is Sparre Andersen’s [28] classical result that, for $|x|\leqslant 1$,
@@
We note that Vysotsky [30] sharpened and generalized Sinaĭ’s persistence probability (1), showing that
${\bf P}(A_{1},\ldots,A_{n}\geqslant 0)\sim Cn^{-1/4},$
More specifically, in [30] it is assumed that the increments of the random walk are $\alpha$-stable, for some $1<\alpha\leqslant 2$, and either right-exponential or right-continuous (see [30] for the precise conditions). Subsequently, Dembo, Ding and Gao [19] extended these results to the case that the increments have finite, positive second moment. See also Gao, Liu and Yang [19] for the case of Gaussian increments.
Finally, we note that continuous analogues of such persistence problems have been studied very recently by Bär, Duraj and Wachtel [19], giving rise to a class of random processes called Kolmogorov diffusions.
@@ 1.3. Sinaĭ excursions
We call $(S_{0},S_{1},\ldots,S_{4n})$ a Sinaĭ excursion if it is a Sinaĭ walk and $S_{4n}=A_{4n}=0$. Such renewal times are only possible at multiples of $4$. We note that Sinaĭ excursions are the discrete analogue of the positive Kolmogorov excursions (from zero and back) studied in [19].
@@
verifying a conjecture in Caravenna and Deuschel [11].
@@
To give a first hint about the connection between $p_{n}$ and $\Xi_{n}$, consider the $2n$ times $t_{1}<\cdots<t_{2n}$ before down steps in a walk $(S_{0},S_{1},\ldots,S_{4n})$ when $S_{t+1}-S_{t}=-1$. By Lemma 3.1 below, if $A_{4n}=0$ then $\sum_{j=1}^{2n}t_{j}=n(4n-1)$. Thus, $\{t_{1},\ldots,t_{2n}\}$ is a subset of $\{1,2,\ldots,4n-1\}$ of size $2n$ that sums to $3n$ mod $4n$. More to the point, as we will see, $\Xi_{n}$ is related to the number of bridges that can be turned into a Sinaĭ excursion, by cyclically shifting their increments.
@@ 1.4. A Sparre Andersen analogue
$\varphi_{n}={\bf P}(A_{1},\ldots,A_{4n}\geqslant 0,\>A_{4n}=S_{4n}=0).$ (6)
By the local limit in [2, Proposition 2.1], it follows that
$n^{2}{\bf P}(A_{4n}=S_{4n}=0)\to\frac{\sqrt{3}}{4\pi}.$
@@ 1.5. Transferring asymptotics
A probability distribution $(q_{n},n\geqslant 0)$ is sub-exponential if $q_{n}/q_{n+1}\to 1$ and its convolution
@@
satisfies that $q_{n}^{*}/q_{n}\to 2$. In [16] it is proved that if $p_{n}$ and $\nu_{n}$ are related by (10), then $\nu_{n}/\lambda$ is sub-exponential if and only if $p_{n}\sim\nu_{n}$ and $\nu_{n}/\nu_{n+1}\to 1$. Intuitively, this follows by the “one big jump principle.” A large value of $L_{1}$ is likely due to one big jump of essentially this value by time $t=1$. As a special case, Hawkes and Jenkins [22] showed that $p_{n}\sim\nu_{n}$ if $\nu_{n}$ is regularly varying with index $\gamma<-1$.
@@ 1.6. Outline
We have shown how Theorem 1.1 follows from Theorem 1.2 and the asymptotics of $\Xi_{n}$. The remainder of the article is devoted to the proof of Theorem 1.2.
@@
Next, we claim that, for integers $10\leqslant m_{1}\leqslant\cdots\leqslant m_{2n}\leqslant 2n-$ 1, we have that $\sum_{i=1}^{2n}m_{i}$ is equal to 00 mod $2n$ if and only if $\sum_{i=1}^{2n}(m_{i}+i)$ is equal to $n$ or $3n$ mod $4n$. Indeed, to see this, simply note that $\sum_{i=1}^{2n}i=n(2n+1)$ is equal to $n$ or $3n$ mod $4n$ (depending on the parity of $n$). This implies that the sub-multisets of $\{1,\dots,2n-1\}$ of size $2n$ that sum to 00 mod $2n$ are in bijection with the subsets of $\{1,\dots,4n-1\}$ of size $2n$ that sum to $n$ or $3n$ mod $n$, and the statement follows. ∎
@@ 3. Times before down steps
denote the sequences of times before down steps, that is, times $10\leqslant t_{1}<\dots<t_{n}\leqslant 2n-$ 1 such that $\Delta B_{t}=-1$.
@@
$\sum_{k=1}^{4n}B_{k}=\sum_{k=0}^{4n-1}(4n-k)\Delta B_{k}=\sum_{k=0}^{4n-1}(4n-k)-2\sum_{j=1}^{2n}(4n-t_{j}),$
where the last step follows from $\Delta B_{k}=1-2\mathbf{1}{\{k\in\{t_{1},\dots,t_{2n}\}\}}$. This simplifies to $-2n(4n-1)+2\sum_{i=1}^{2n}t_{i}$, as claimed. ∎
@@ 4. Sparre Andersen for Sinaĭ excursions
For this, we use the observation that $\Phi_{n}$ is a renewal sequence. A renewal sequence enumerates structures that can be decomposed into a series of irreducible parts. Formally, $A_{n}$ is a renewal sequence if its generating function $A(x)=\sum_{n=0}^{\infty}A_{n}x^{n}$ can be expressed as
$A(x)=\sum_{m=0}^{\infty}[A^{(1)}(x)]^{m}=\frac{1}{1-A^{(1)}(x)}$
where $A^{(1)}(x)=\sum_{n=1}^{\infty}A_{n}^{(1)}x^{n}$ is the generating function for the number $A_{n}^{(1)}$ of irreducible structures of length $n$. See, e.g., Feller [18] for details.
We see that $\Phi_{n}$ is a renewal sequence because any Sinaĭ excursion can be decomposed into a series of irreducible Sinaĭ excursions whose area process only takes the value 00 at the first and the last step. More specifically, let $\Phi_{n}^{(1)}$ be the number of walks $(S_{0},S_{1},\ldots,S_{4n})$ for which $S_{0}=A_{0}=0$, $S_{4n}=A_{4n}=0$ and $A_{1},\ldots,A_{4n-1}>0$. Then, the generating function $\Phi(x)=\sum_{n}\Phi_{n}x^{n}$ of $\Phi_{n}$ can be expressed as
$\Phi(x)=\frac{1}{1-\Phi^{(1)}(x)},$
where $\Phi^{(1)}(x)=\sum_{n}\Phi_{n}^{(1)}x^{n}$ is the generating function for the number $\Phi^{(1)}_{n}$ of irreducible Sinaĭ excursions.
Then, the key to proving (13) is the following lemma, proved by Bassan and the authors in [7].
Lemma 4.1 (Lemma 2 (1) in [7]).
@@
where $A_{n}^{\prime}$ is the number of pairs $(X,s)$, where $X\in\mathcal{A}_{n}$ and $0\leqslant s<\ell$ is an integer, where $\ell$ is the size of the first irreducible part of $X$.
Its proof is simple, so we include it here.
Proof.
Note that $nA_{n}$ enumerates the pairs $(X,i)$, where $X\in\mathcal{A}_{n}$ and $0\leqslant i<n$ is an integer. We call such pairs marked objects of size $n$. For a marked object $(X,i)$ of size $n$, consider the sub-object consisting of the irreducible part containing $i$ and all subsequent parts. This sub-object is of size $k$, for some $1\leqslant k\leqslant n$, and has a mark somewhere in its first irreducible part. All other previous parts form an (unmarked) object of size $n-k$. This implies that there are $\sum_{k=1}^{n}A_{k}^{\prime}A_{n-k}$ many such pairs $(X,i)$. See Figure 2. ∎
Figure 2. Marked objects of size $n$ ($nA_{n}$ many) can be split into an object of size $n-k$ ($A_{n-k}$ many) and an object of size $k$ with a mark in its first irreducible part ($A_{k}^{\prime}$ many). Vertical lines delimit irreducible parts. Recurrence (14) follows, summing over $k$.
Lemma 4.1 and the observation that $\Phi_{n}$ is a renewal sequence imply the following lemma (recalling that renewal times when $S_{4k}=A_{4k}=0$ are only possible at multiples of $4$).
Lemma 4.2.
Let $\Phi^{\prime}_{n}$ denote the number of pairs $({\mathcal{B}},s)$, where ${\mathcal{B}}$ is a Sinaĭ excursion of length $4n$ and $0\leqslant s<\ell$ is an integer, where $\ell$ the length of the first irreducible Sinaĭ excursion of ${\mathcal{B}}$. Then,
$n\Phi_{n}=\sum_{k=1}^{n}\Phi_{k}^{\prime}\Phi_{n-k}.$
Finally, to prove (13) and thereby complete the proof of Theorem 1.2, we show the following.
@@
  • from the set of all pairs $({\mathcal{B}},j)$, where ${\mathcal{B}}$ is a Sinaĭ excursion of length $4n$, with first positive Sinaĭ excursion of length $\ell=4k$, and $1\leqslant j\leqslant 4k$ is an integer,
@@
To describe $\Upsilon$, consider ${\mathcal{B}}$ as above. Let
@@
be the times before up steps in the first positive Sinaĭ excursion of ${\mathcal{B}}$. Let ${\mathcal{B}}^{(j)}$ be the bridge obtained from ${\mathcal{B}}$ by cyclically shifting ${\mathcal{B}}$ to the left by $i_{j}$. In other words, the $k$th increment of ${\mathcal{B}}^{(j)}$ is the $(k+i_{j})$th (understood mod $4n$) increment of ${\mathcal{B}}$. In particular, ${\mathcal{B}}^{(1)}={\mathcal{B}}$. We let $\Upsilon({\mathcal{B}},j)$ to be the set of times before down steps in ${\mathcal{B}}^{(j)}$. Lemma 3.1 implies that, because ${\mathcal{B}}$ is a Sinaĭ excursion, its times before down steps sum to $n(4n-1)\equiv 3n$ mod $4n$. We shift all of these $2n$ times by the same amount to obtain the times before down steps in ${\mathcal{B}}^{(j)}$, so these sum to $n$ or $3n$ mod $4n$. See Figure 3 for an example.
@@
5. A related application
Finally, let us conclude with another, related application of our current techniques.
As a consequence of considerably more technical arguments than those in the current article, we [15, Corollary 4] recently proved that
@@ 5. A related application
where $\tau=\inf\{t:S_{t}=0,\>A_{t}\leqslant 0\}$. Informally, $\tau$ is the first time that a random walk is “at risk” of not being Sinaĭ (i.e., if $A_{\tau}<0$). Theorem 1.1 above implies that ${\bf P}(A_{\tau}=0)=1-e^{-\lambda}$, where $\lambda$ is as in (8), so that
@@ 6. Acknowledgments
SD acknowledges the financial support of the CogniGron research center and the Ubbo Emmius Funds (University of Groningen). BK was partially supported by a Florence Nightingale Bicentennial Fellowship (Oxford Statistics) and a Senior Demyship (Magdalen College). BK gratefully acknowledges Vasu Tewari for indicating the work of von Sterneck.